# Semiparametric Bayesian Difference-in-Differences

### v1

@@ 4.1 Assumptions for Inference
where the conditional density $f_{\Delta Y|D,X}$ is a function of the conditional mean $m$ by the exponential family restriction given in (2.4). We write $\mathcal{L}_{\Pi}(\sqrt{n}(\tau_{\eta}-\widehat{\tau})|Z^{(n)})$ for the marginal posterior law of $\sqrt{n}(\tau_{\eta}-\widehat{\tau})$.
@@
The posterior contraction rate for the conditional mean can be derived by modifying the classical results of Ghosal, Ghosh, and van der Vaart (2000). In the semiparametric Bayesian literature, the requirement $\varepsilon_{n}=o(n^{-1/4})$ is stated explicity in order to kill some second-order remainder terms; see Condition (C) in Castillo (2012). This also agrees with the usual cut-off rate of the nonparametric components in frequentist semiparametric models (Newey, 1994). Note that the ATT $\tau_{\eta}$ is linear in $m_{\eta}$, so that we do not need to deal with these second-order terms. Nevertheless, the posterior contraction rate also play a role in the next two assumptions related to the stochastic equicontinuity and prior stability. For the concrete example involving the Hölder class for the conditional mean function, we need to impose a sufficient smoothness condition, so that this contraction rate indeed satisfy $\sqrt{n}\varepsilon^{2}_{n}=o(1)$, i.e., $\varepsilon_{n}=o(n^{-1/4})$.
@@ 6 Illustration using Squared Exponential Process Priors
Proposition 6.2 requires $\sqrt{s_{\pi}\,s_{m}}>p/2$, which represents a trade-off between the smoothness requirement for $m_{0}$ and $\pi_{0}$. This encapsulate the double robustness; i.e., a lack of smoothness of the conditional mean function $m_{0}$ can be mitigated by exploiting the regularity of the propensity score and vice versa.
@@ 7.1 Simulation Evidence
Our nonparametric Bayesian (hereinafter Bayes) and the double robust Bayesian (DR Bayes) methods are implemented following Algorithms 1 and 2 in Section 3, using the MATLAB package $\mathtt{GPML}$ to draw posteriors. Both Bayesian methods are implemented based on $5000$ posterior draws. For DR Bayes, the propensity score is estimated by logistic regression, and the pilot conditional mean estimator $\widehat{m}$, used for posterior correction, is computed as the posterior mean of $m^{s}_{\eta}$ using uncorrected Gaussian process priors. The tuning parameter $\varsigma_{n}$, corresponding to the standard deviation of the adjusted prior, is set as $\nu\log n_{0}/(\sqrt{n_{c}}\,\Gamma_{n})$, where $\Gamma_{n}=\sum_{i=1}^{n}|\widehat{\gamma}(0,X_{i})(1-D_{i})|/n_{c}$, where $n_{c}$ denotes the number of units in the control group. In Supplementary Appendix E, Table A1 demonstrates that the performance of DR Bayes is stable with respect to the value of $\varsigma_{n}$. DR Bayes in Table 1 uses the full sample twice in computing the prior/posterior adjustments and the posteriors of the conditional mean function. As shown in Table A2 in Supplementary Appendix E, results from sample splitting are comparable to those in Table 1.
We also compare the Bayesian methods to several frequentist DiD estimators. DR corresponds to the improved doubly robust DiD estimator proposed by Sant’Anna and Zhao (2020). OR, the outcome regression approach, refers to the sample analog of (2.1) where the conditional mean $m_{0}$ is estimated by a linear regression of $\Delta Y_{i}$ on $X_{i}$ using the sample of the control arm. IPW refers to the propensity score reweighed estimator (Abadie, 2005), and TWFE corresponds to the standard two-way-fixed effect model that regresses $Y_{it}$ on $D_{i}$, $t$, the interaction $D_{i}\times t$ and $X_{i}$. DML corresponds to the double/debiased machine learning estimator of Chernozhukov, Chetverikov, Demirer, Duflo, Hansen, and Newey (2017), where the nuisance function $\pi_{0}$ is estimated by logistic LASSO and $m_{0}$ estimated by random forests. [^4] We apply random forest to estimate $m_{0}$ to cope with the nonlinear function forms in Designs II and IV, and to match DML with our Bayesian procedures that estimate $m_{0}$ nonparametrically. Frequentist DiD estimators, except for DML, are implemented using the R package $\mathtt{DRDID}$, while DML is implemented using the R package $\mathtt{DoubleML}$. Table 1 presents the finite sample (mean) bias of the point estimator, coverage probability (CP) and the average length (CIL) of the $95\%$ credible/confidence interval for the Bayesian and frequentist methods mentioned above.
@@ 7.2 Empirical illustration: Minimum Wage
As we see from Table 3, nonparametric Bayes and DR Bayes yield ATT point estimates ranging from $1.91$ to $2.03$ and confidence intervals covering 00 with the length from $5.53$ to $6.68$. The frequentist estimators, especially the doubly robust ones (DR and DML), yield longer confidence intervals. If we do not trim the propensity score, then the failure of the overlap condition prevents us from using the estimators that involve inverse propensity score. Among estimators that do not use the propensity score, nonparametric Bayes gives an ATT estimate of $1.935$, with a 95% confidence interval of $[-0.460,4.341]$, for the full sample without any trimming ($t=0,\bar{n}_{1}=249,\bar{n}_{0}=58$). OR yields an estimated ATT of $3.351$, with a 95% CI of $[-1.233,7.936]$. TWFE provides an estimated ATT of $2.635$, with a 95% CI of $[-0.622,5.891]$. It turns out that the nonparametric Bayesian method continue to yield stable results when the overlap condition is in doubt. In sum, our Bayesian methods, which allows a flexible form of the conditional mean function $m_{0}(x)$ as well as a rich set of covariate, generate comparable ATT estimate with the original findings in Card and Krueger (1994). [^6] When covariates are not included in the model, Card and Krueger (1994) report the difference-in-difference estimate of $2.76$ (standard error $1.36$), and the regression adjusted model with controls for chain and ownership dummies yields an estimate of $2.30$ (standard error $1.20$). Therefore, our Bayesian DiD methods confirms the robustness of findings in the classic literature against model specifications.
@@ 8.1 Extension to Multiple Periods and Staggered Entry
Suppose the available panel data consists of $T$ periods indexed by $t=1,\dots,T$ and the earliest treatment intervention occurs at period $S$. We assume that the treatment intervention remains once a unit get treated. As a result, the entire path of treatment assignment for each unit can be summarized by his/her first treated period (cohort), denoted by the cohort variable $G_{i}\in\{S,\dots,T,\infty\}$, where $G_{i}=\infty$ means the unit $i$ never gets treated. Let the cohort indicators $D_{ig}$ denote whether unit $i$ first receives treatment in period $g\in{S,\ldots,T,\infty}$, where $D_{i\infty}=1$ indicates that unit $i$ never receives treatment.
@@
Suppose a vector of pre-treatment covariates $X_{i}$ is also available, a vector of dimension $p$, with the distribution $F_{0}$ and the density $f_{0}$. The researcher observes independent and identically distributed observations of $(Y_{i1},\dots,Y_{iT},D_{iS},\dots,D_{iT},D_{i\infty},X_{i})$.
@@ 9 Conclusion
This paper introduces new semiparametric Bayesian procedures that satisfy the Bernstein-von Mises Theorem. Our first proposal, based on standard Gaussian process priors, provides a Bayesian analog to the outcome regression in Heckman, Ichimura, and Todd (1997). In simulations, we show that it performs well in models that are not overly complex, does not require a propensity score specification, and adapts to the functional form of the conditional mean function. Our second proposal possessing double rubustness is more suited to complex models. Here, we observe that it provides useful complements to existing frequentist double-robust methods. While our focus is primarily on the Difference-in-Differences (DiD) panel data case, we discuss extensions to the repeated cross-section case and staggered interventions.
@@ Appendix A Proofs of Main Results
$\upsilon_{\eta}t\sqrt{n}(\tau_{\eta}-\widehat{\tau})-tR_{n,\eta}$ $=\underbrace{\frac{t}{\sqrt{n}}\sum_{i}\Big{(}\mathbb{E}_{\eta}\left[\frac{D}{\pi_{0}}(\Delta Y-m_{\eta}(X)-\tau_{0})\right]-\frac{D_{i}}{\pi_{0}}(\Delta Y_{i}-m_{\eta}(X_{i})+\tau_{0})\Big{)}}_{=t\sqrt{n}(P_{\eta}-\mathbb{P}_{n})\left[\frac{d}{\pi_{0}}(\Delta y-m_{\eta}(x)-\tau_{0})\right]}$
@@
In this supplement, $C>0$ denotes a generic constant, whose value might change line by line. We introduce additional subscripts when there are multiple constant terms in the same display. For two sequences $a_{n},b_{n}$, we write $a_{n}\lesssim b_{n}$, if $a_{n}\leq Cb_{n}$.
@@ C.2 Expansions
The next lemma is about smaller order terms in the proof of our BvM theorems. Note that the posterior law of $F_{\eta}$ coincides with the Bayesian bootstrap. Here, the negligibility of those terms refers to the randomness with respect to the Bayesian bootstrap weights (for which we use $P_{M}$ to highlight this dependence), conditional on the observed data. We refer readers to Page 2891 in cheng2010bootstrap for comprehensive discussion about disentangling the sources of randomness coming from the observed data and the Bayesian bootstrap weights. Formally, we define $\Delta_{n}=o_{P_{M}}(1)$ in $P_{0}$-probability, if for any small positive $\epsilon$ and $\delta$, it holds $P_{0}\left(P_{M|Z^{(n)}}(|\Delta_{n}|>\epsilon)>\delta\right)\to 0$ In addition, we define $\Delta_{n}=O_{P_{M}}(1)$ in $P_{0}$-probability, if for any small positive $\delta$, there exists a large enough $C$ such that $P_{0}\left(P_{M|Z^{(n)}}(|\Delta_{n}|>C)>\delta\right)\to 0$. For the next result, recall the definition he remainder term $R_{n,\eta}$ given in the proof of Theorem 4.1 by
@@
By the moment condition for the envelope function in Assumption 3 (i), The first term is $O_{P_{0}}(1)$ and the second term is $O_{P_{M}}(1)$ in $P_{0}$-probability. By the relationship in (71) of cheng2010bootstrap, the remainder term $R_{n,\eta}$ converges to zero in $P_{Z}$-probability, conditional on the data.

### v2

@@ 4.1 Assumptions for Inference
where the conditional density $f_{\Delta Y|D,X}$ is a function of the conditional mean $m$ by the exponential family restriction given in (2.4). Here, we used the fact that independent priors are placed on the conditional mean $m$ and the distribution function $F$. We write $\mathcal{L}_{\Pi}(\sqrt{n}(\tau_{\eta}-\widehat{\tau})|Z^{(n)})$ for the marginal posterior law of $\sqrt{n}(\tau_{\eta}-\widehat{\tau})$.
@@
The posterior contraction rate for the conditional mean can be derived by modifying the classical results of Ghosal, Ghosh, and van der Vaart (2000). In the related literature, the requirement $\varepsilon_{n}=o(n^{-1/4})$ is stated explicitly in order to eliminate second-order remainder terms; see Condition (C) in Castillo (2012). This also aligns with the usual cut-off rate of the nonparametric components in frequentist semiparametric models (Newey, 1994). Note that the ATT $\tau_{\eta}$ is linear in $m_{\eta}$, so that we do not need to deal with these second-order terms. Nevertheless, the posterior contraction rate also plays a crucial role in the next two assumptions related to the stochastic equicontinuity and prior stability. For the concrete example involving the Hölder class for the conditional mean function, we need to impose sufficient smoothness so that this contraction rate indeed satisfies $\sqrt{n}\varepsilon^{2}_{n}=o(1)$. If the exponential family structure (2.4) is misspecified, the posterior contracts to the point in the support of the prior that is closest to the true distribution (as measured by the Kullback-Leibler divergence). Specifically, if one starts with the Gaussian model, the contraction result required by Assumption 2 can be established utilizing Theorem 4.1 of Kleijn and van der Vaart (2006), provided that the true conditional mean lies within the prior’s support.
@@ 6 Illustration using Squared Exponential Process Priors
[…]
@@ 7.1 Simulation Evidence
Our nonparametric Bayesian (hereinafter Bayes) and the double robust Bayesian (DR Bayes) methods are implemented following Algorithms 1 and 2 in Section 3, using the MATLAB package $\mathtt{GPML}$ to draw posteriors. Both Bayesian methods are implemented based on $5000$ posterior draws. Here, we apply the exponential family specification in (2.4) to the Gaussian case. The resulting posterior distribution of the conditional mean function is available in closed form (see Supplementary Appendix D for details), eliminating the need for computationally costly Monte Carlo samplers like MCMC. [^4] If the conditional density function in (2.4) belongs to another distribution in the exponential family, the posterior of the conditional mean can also be approximated using an analytical approximation, such as the Laplace method, see Riihimäki and Vehtari (2014). For DR Bayes, the propensity score is estimated by logistic regression, and the pilot conditional mean estimator $\widehat{m}$, used for posterior correction, is computed as the posterior mean of $m^{s}_{\eta}$ using uncorrected Gaussian process priors. The tuning parameter $\varsigma_{n}$, corresponding to the standard deviation of the adjusted prior, is set as $\nu\log n_{0}/(\sqrt{n_{c}}\,\Gamma_{n})$, where $\Gamma_{n}=\sum_{i=1}^{n}|\widehat{\gamma}(0,X_{i})(1-D_{i})|/n_{c}$, where $n_{c}$ denotes the number of units in the control group. In Supplementary Appendix E, Table A1 demonstrates that the performance of DR Bayes is stable with respect to the value of $\varsigma_{n}$. DR Bayes in Table 1 uses the full sample twice in computing the prior/posterior adjustments and the posteriors of the conditional mean function. As shown in Table A2 in Supplementary Appendix E, results from sample splitting are comparable to those in Table 1.
We also compare the Bayesian methods to several frequentist DiD estimators. DR corresponds to the improved doubly robust DiD estimator proposed by Sant’Anna and Zhao (2020). OR, the outcome regression approach, refers to the sample analog of (2.1) where the conditional mean $m_{0}$ is estimated by a linear regression of $\Delta Y_{i}$ on $X_{i}$ using the sample of the control arm. IPW refers to the propensity score reweighed estimator (Abadie, 2005), and TWFE corresponds to the standard two-way-fixed effect model that regresses $Y_{it}$ on $D_{i}$, $t$, the interaction $D_{i}\times t$ and $X_{i}$. DML corresponds to the double/debiased machine learning ATT estimator of Chernozhukov, Chetverikov, Demirer, Duflo, Hansen, and Newey (2017) or Chang (2020), where the nuisance function $\pi_{0}$ is estimated by logistic LASSO and $m_{0}$ estimated by random forests. [^5] We apply random forest to estimate $m_{0}$ to cope with the nonlinear function forms in Designs II and IV, and to match DML with our Bayesian procedures that estimate $m_{0}$ nonparametrically. Frequentist DiD estimators, except for DML, are implemented using the R package $\mathtt{DRDID}$, while DML is implemented using the R package $\mathtt{DoubleML}$. Table 1 presents the finite sample (mean) bias of the point estimator, coverage probability (CP) and the average length (CIL) of the $95\%$ credible/confidence interval for the Bayesian and frequentist methods mentioned above.
@@ 7.2 Empirical illustration: Minimum Wage
As we see from Table 3, all methods produce positive but insignificant ATT estimates for the impact of minimum wage, which is in line with the findings of Card and Krueger (1994). For example, nonparametric Bayes and DR Bayes yield ATT point estimates ranging from $1.907$ to $2.024$ and confidence intervals covering 00 with the length from $5.514$ to $6.683$. Bayesian methods also provide shorter confidence intervals than frequentist methods including the widely-used TWFE estimator.
If we do not trim the propensity score, then the failure of the overlap condition prevents us from using the estimators that involve inverse propensity score. Among estimators that do not use the propensity score, nonparametric Bayes gives an ATT estimate of $1.935$, with a 95% confidence interval of $[-0.460,4.341]$, for the full sample without any trimming ($t=0,\bar{n}_{1}=249,\bar{n}_{0}=58$). OR yields an estimated ATT of $3.351$, with a 95% CI of $[-1.233,7.936]$. TWFE provides an estimated ATT of $2.635$, with a 95% CI of $[-0.622,5.891]$. It turns out that our semiparametric Bayesian method continue to yield stable results when the overlap condition is in doubt. In sum, our Bayesian methods, which allows a flexible form of the conditional mean function $m_{0}(x)$ as well as a rich set of covariate, generate comparable ATT estimate with the original findings in Card and Krueger (1994). [^7] When covariates are not included in the model, Card and Krueger (1994) report the difference-in-difference estimate of $2.76$ (standard error $1.36$), and the regression adjusted model with controls for chain and ownership dummies yields an estimate of $2.30$ (standard error $1.20$). Therefore, our Bayesian DiD methods confirms the robustness of findings in the classic literature against model specifications.
@@ 8.1 Extension to Multiple Periods and Staggered Entry
Suppose the available panel data consists of $T$ periods indexed by $t=1,\dots,T$ and the earliest treatment intervention occurs at period $S$. We assume that the treatment intervention remains once a unit get treated. As a result, the entire path of treatment assignment for each unit can be summarized by his/her first treated period (cohort), denoted by the cohort variable $G_{i}\in\{S,\dots,T,\infty\}$, where $G_{i}=\infty$ means the unit $i$ never gets treated. Let the cohort indicators $D_{ig}$ denote whether unit $i$ first receives treatment in period $g\in\{S,\ldots,T,\infty\}$, where $D_{i\infty}=1$ indicates that unit $i$ never receives treatment.
@@
Suppose a vector of pre-treatment covariates $X_{i}$ is also available, a vector of dimension $p$, with the distribution $F_{0}$ and the density $f_{0}$. The researcher observes independent and identically distributed observations of $(Y_{i1},\dots,Y_{iT},D_{iS},\dots,D_{iT},D_{i\infty},X_{i})$, $i=1,\ldots,n$.
@@
Corollary 8.1 pertains to inference on cohort-specific ATTs and establishes BvM results for our two Bayesian methods, employing either standard Gaussian process priors or prior/posterior adjustments via the cohort-specific propensity score. We note that extending this framework to aggregate ATTs is highly non-trivial, as it requires a joint modeling of outcome variables across different cohorts and time periods. Hence, distinct prior and likelihood specifications in the Bayesian methodology, as well as prior/posterior adjustments of the double robust version, are needed. A thorough investigation is therefore left for future research.
@@ 9 Conclusion
This paper introduces new semiparametric Bayesian procedures that satisfy the Bernstein-von Mises results in the DiD setup. Our first proposal, based on standard Gaussian process priors, provides a Bayesian analog to the outcome regression in Heckman, Ichimura, and Todd (1997). Through simulations, we show that it performs well in models that are not overly complex and, since no propensity score specification is required, it is not sensitive to overlap issues. Our second, double robust proposal incorporates prior/posterior corrections based on estimated propensity scores. In simulations it works well for complex models, i.e., when the number of covariates is large. Overall, our Bayesian methods exhibit remarkable finite sample performance, while adapting to the functional form of the conditional mean function. Although our focus is primarily on the DiD panel data case, we also discuss extensions to the repeated cross-section case and staggered interventions.
@@ Appendix A Proofs of Main Results
$\upsilon_{\eta}t\sqrt{n}(\tau_{\eta}-\widehat{\tau})-tR_{n,\eta}$ $=\underbrace{\frac{t}{\sqrt{n}}\sum_{i}\left(\mathbb{E}_{\eta}\left[\frac{D}{\pi_{0}}(\Delta Y-m_{\eta}(X)-\tau_{0})\right]-\frac{D_{i}}{\pi_{0}}(\Delta Y_{i}-m_{\eta}(X_{i})+\tau_{0})\right)}_{=t\sqrt{n}(P_{\eta}-\mathbb{P}_{n})\left[\frac{d}{\pi_{0}}(\Delta y-m_{\eta}(x)-\tau_{0})\right]}$
@@
In this supplement, $C>0$ denotes a generic constant, whose value might change line by line. We introduce additional subscripts when there are multiple constant terms in the same display.
@@ C.2 Expansions
The next lemma is about smaller order terms in the proof of our BvM theorems. Note that the posterior law of $F_{\eta}$ coincides with the Bayesian bootstrap. Here, the negligibility of those terms refers to the randomness with respect to the Bayesian bootstrap weights (for which we use $P_{M}$ to highlight this dependence), conditional on the observed data. We refer readers to Page 2891 in Cheng and Huang (2010) for comprehensive discussion about disentangling the sources of randomness coming from the observed data and the Bayesian bootstrap weights. Formally, we define $\Delta_{n}=o_{P_{M}}(1)$ in $P_{0}$-probability, if for any small positive $\epsilon$ and $\delta$, it holds $P_{0}\left(P_{M|Z^{(n)}}(|\Delta_{n}|>\epsilon)>\delta\right)\to 0$ In addition, we define $\Delta_{n}=O_{P_{M}}(1)$ in $P_{0}$-probability, if for any small positive $\delta$, there exists a large enough $C$ such that $P_{0}\left(P_{M|Z^{(n)}}(|\Delta_{n}|>C)>\delta\right)\to 0$. For the next result, recall the definition he remainder term $R_{n,\eta}$ given in the proof of Theorem 4.1 by
@@
By the moment condition for the envelope function in Assumption 3 (i), The first term is $O_{P_{0}}(1)$ and the second term is $O_{P_{M}}(1)$ in $P_{0}$-probability. By the relationship in (71) of Cheng and Huang (2010), the remainder term $R_{n,\eta}$ converges to zero in $P_{Z}$-probability, conditional on the data.